OpenStudy (anonymous):

Solve log3 x – 1 = 3

OpenStudy (anonymous):

is this th problem: \(\large log_3(x-1)=3 \)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if yes, then all you need to do is write this log equation in exponential form. that is, use this fact: \[\huge log_bM=x\Leftrightarrow b^x=M\]

OpenStudy (anonymous):

i dont get

OpenStudy (anonymous):

\[\huge log_3(x-1)=3\Leftrightarrow 3^3=(x-1) \] now can you solve it?

OpenStudy (anonymous):

ooh ok :)

OpenStudy (anonymous):

x=28

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

thanks